# Ordered Momentum for Asynchronous SGD

### v2

@@ C.3.2 Notation
We define $\hat{\tau}_{k,t}=t-ite(k,t)$, where $k\in[K]$ and $t\in[T+1]$. $\hat{\tau}_{k,t}$ denotes the current delay of the gradient on worker $k$ at iteration $t$, which is the number of iterations that have happened since $ite(k,t)$. It’s easy to verify that $\hat{\tau}_{k_{t},t}=t-ite(k_{t},t)=\tau_{t}$, where $t\in[T]$.
where $k\in[K]$, $t\in[T]$. It’s easy to verify that $ite(k,next(k,t))=ite(k,t),\forall k\in[K],t\in[T]$.
@@
We also define an auxiliary sequence $\hat{\eta}_{k,t}$ for the learning rate. $\hat{\eta}_{k,t}$ denotes the learning rate corresponding to the gradient on worker $k$ at iteration $t$.
@@
where $k\in[K]$ and $t\in[T]$.
@@
It’s easy to verify that $\hat{\eta}_{k_{t},t}=\eta_{next(k_{t},t)}=\eta_{t}=\eta_{next(k_{t},ite(k_{t},t))}=\hat{\eta}_{k_{t},ite(k_{t},t)}$, where $t\in[T]$.
@@ C.3.3 Convergence Analysis for OrMo-DA
$~{}~{}~{}~{}+\frac{L^{2}\hat{\eta}_{k_{t-1},t}}{1-\beta}\left\|\hat{{\bf y}}_{t}-\hat{{\bf w}}_{t}\right\|^{2}+\frac{L^{2}\hat{\eta}_{k_{t-1},t}}{1-\beta}\left\|\hat{{\bf w}}_{t}-{\bf w}_{t}\right\|^{2}$
@@
$\mathbb{E}\left\|\hat{{\bf u}}_{t}\right\|^{2}$ $\leq 2\mathbb{E}\left\|\beta^{\lfloor\frac{t+K-2}{K}\rfloor}\left(\sum_{k\in[K]}\hat{\eta}_{k,0}\nabla F({\bf w}_{0})\right)+\sum_{s=1}^{\lfloor\frac{t+K-2}{K}\rfloor}\beta^{\lfloor\frac{t+K-2}{K}\rfloor-s}\left(\sum_{j=(s-1)K+1}^{\min\{sK,t-1\}}\hat{\eta}_{k_{j-1},j}\nabla F({\bf w}_{j})\right)\right\|^{2}$
@@
$\leq\frac{2K}{1-\beta}\left[\sum_{k\in[K]}\beta^{\lfloor\frac{t+K-2}{K}\rfloor}{\left(\hat{\eta}_{k,0}\right)}^{2}\left\|\nabla F({\bf w}_{0})\right\|^{2}\right.$
@@
$\sum_{t=1}^{T-1}\mathbb{E}\left(\hat{\eta}_{k_{t-1},t}\left\|\hat{{\bf w}}_{t}-{\bf w}_{t}\right\|^{2}\right)$ $\leq\frac{\eta K}{2L{(1-\beta)}^{2}}\left[\sum_{t=1}^{T-1}\hat{\eta}_{k_{t-1},t}\left\|\nabla F({\bf w}_{t})\right\|^{2}+\left(\sum_{k\in[K]}\hat{\eta}_{k,0}\right)\left\|\nabla F({\bf w}_{0})\right\|^{2}\right]$
$~{}~{}~{}~{}+\frac{\eta\sigma^{2}}{2L{(1-\beta)}^{2}}\left(\sum_{t=1}^{T-1}\hat{\eta}_{k_{t-1},t}+\sum_{k\in[K]}\hat{\eta}_{k,0}\right)$
@@
$=\frac{2\eta}{{(1-\beta)}^{2}}\sum_{j=1}^{T-2}{\left(\hat{\eta}_{k_{j-1},j}\right)}^{2}\left(\sigma^{2}+K\mathbb{E}\left\|\nabla F({\bf w}_{j})\right\|^{2}\right)\left(next(k_{j-1},j)-j\right)$
@@
$\eta_{t}=\left\{\begin{aligned} &\eta&\tau_{t}\leq 2K\\ &\min\{\eta,\frac{1}{4L\tau_{t}}\}&\tau_{t}>2K,\end{aligned}\right.$
@@
$\mathbb{E}\left\|\nabla F(\tilde{{\bf w}}_{T})\right\|^{2}\leq\mathcal{O}(\sqrt{\frac{L\sigma^{2}}{T}}+\frac{KL}{T}).$
where $\tilde{{\bf w}}_{T}$ is randomly chosen from ${\bf w}_{0},{\bf w}_{1},\cdots,{\bf w}_{T-1}$ according to (17).
@@
$\leq F({\bf w}_{0})+\left[\frac{L{\left(\sum_{k\in[K]}\hat{\eta}_{k,0}\right)}^{2}}{2{(1-\beta)}^{2}}-\frac{\sum_{k\in[K]}\hat{\eta}_{k,0}}{1-\beta}\right]\left\|\nabla F({\bf w}_{0})\right\|^{2}$
$~{}~{}~{}~{}+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{k\in[K]}{\left(\hat{\eta}_{k,0}\right)}^{2}$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T-1}\left(\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\left\|\nabla F({\bf w}_{t})\right\|^{2}\right)+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}\right]$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T-1}\left(\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\left\|\nabla F({\bf w}_{t})\right\|^{2}\right)+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}\right]$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T-1}\left(\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\left\|\nabla F({\bf w}_{t})\right\|^{2}\right)+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}\right]$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T-1}\left(\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\left\|\nabla F({\bf w}_{t})\right\|^{2}\right)+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}\right]$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T-1}\left(\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\left\|\nabla F({\bf w}_{t})\right\|^{2}\right)+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}\right]$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T-1}\left(\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\left\|\nabla F({\bf w}_{t})\right\|^{2}\right)+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}\right]$
@@
$\leq\sum_{t=1}^{T-1}\mathbb{E}\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{4(1-\beta)}\right)\left\|\nabla F({\bf w}_{t})\right\|^{2}$
@@
$\leq-\frac{1}{8\left(1-\beta\right)}\mathbb{E}\left(\sum_{t=1}^{T-1}\hat{\eta}_{k_{t-1},t}\left\|\nabla F({\bf w}_{t})\right\|^{2}+\sum_{k\in[K]}\hat{\eta}_{k,0}\left\|\nabla F({\bf w}_{0})\right\|^{2}\right)$
@@
$\frac{1}{\sum_{t=0}^{T-1}\tilde{\eta}_{t}}\mathbb{E}\left(\sum_{t=0}^{T-1}\tilde{\eta}_{t}\left\|\nabla F({\bf w}_{t})\right\|^{2}\right)\leq\frac{8\left(1-\beta\right)\left(F({\bf w}_{0})-F^{*}\right)}{\sum_{t=0}^{T-1}\tilde{\eta}_{t}}+\frac{16\eta L\sigma^{2}}{{(1-\beta)}^{2}},$
where $\tilde{\eta}_{0}=\sum_{k\in[K]}\hat{\eta}_{k,0}$ and $\tilde{\eta}_{t}=\hat{\eta}_{k_{t-1},t}(t\geq 1)$.
Then, we analyse the lower bound of $\sum_{t=0}^{T-1}\tilde{\eta}_{t}$.
@@
$\hat{\tau}_{k,t+1}=\left\{\begin{aligned} &\hat{\tau}_{k,t}+1&k\neq k_{t},\\ &0=\hat{\tau}_{k,t}-\tau_{t}&k=k_{t}.\end{aligned}\right.$
Since $\sum_{k\in[K]}\hat{\tau}_{k,0}=0$, we have $\sum_{k\in[K]}\hat{\tau}_{k,T}+\sum_{t=0}^{T-1}\tau_{t}=(K-1)T$. Moreover, $\hat{\tau}_{k_{T-1},T}=T-ite(k_{T-1},T)=0$. We get that $\sum_{k\in[K],k\neq k_{T-1}}\hat{\tau}_{k,T}+\sum_{t=0}^{T-1}\tau_{t}=(K-1)T$. Thus, at least $\frac{T}{2}$ delays are smaller than $2K$. $\sum_{t=0}^{T}\tilde{\eta}_{t}=\sum_{t=1}^{T-1}\hat{\eta}_{k_{t-1},t}+\sum_{k\in[K]}\hat{\eta}_{k,0}\geq\frac{T\eta}{2}$.
$\frac{1}{\sum_{t=0}^{T-1}\tilde{\eta}_{t}}\sum_{t=0}^{T-1}\tilde{\eta}_{t}\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}\leq\frac{16(1-\beta)}{T\eta}\left(F(\hat{{\bf w}}_{0})-F^{*}\right)+\frac{16\eta L\sigma^{2}}{{(1-\beta)}^{2}}.$
If we choose an output $\tilde{{\bf w}}_{T}$ according to
${\mathbb{P}}(\tilde{{\bf w}}_{T}={\bf w}_{t})\propto\tilde{\eta}_{t},$ (17)
where $t\in[T]$ and let $\eta=\min\{\frac{(1-\beta)^{2}}{8KL},\sqrt{\frac{(1-\beta)^{3}\left(F(\hat{{\bf w}}_{0})-F^{*}\right)}{TL\sigma^{2}}}\}$, we have that
$\mathbb{E}\left\|\nabla F(\tilde{{\bf w}}_{T})\right\|^{2}\leq\mathcal{O}\left(\sqrt{\frac{L\sigma^{2}}{T}}+\frac{KL}{T}\right).$

### v3

@@ C.3.2 Notation
where $k\in[K]$, $t\in[T]$. It’s easy to verify that $ite(k,next(k,t))=ite(k,t),next(k,ite(k,t))=next(k,t)$, where $k\in[K],t\in[T]$.
We define $\hat{\tau}_{k,t}=t-ite(k,t)$, where $k\in[K]$ and $t\in[T+1]$. $\hat{\tau}_{k,t}$ denotes the current delay of the gradient ${\bf g}_{ite(k,t)}^{k}$ at iteration $t$, which is the number of iterations that have happened since $ite(k,t)$. It’s easy to verify that $\hat{\tau}_{k_{t},t}=t-ite(k_{t},t)=\tau_{t}$, where $t\in[T]$.
We also define an auxiliary sequence $\hat{\eta}_{k,t}$ for the adaptive learning rates. $\hat{\eta}_{k,t}$ denotes the learning rate corresponding to the gradient ${\bf g}_{ite(k,t)}^{k}$.
@@
where $k\in[K]$ and $t\in[T]$. Since $next(k,t)=next(k,ite(k,t))$, we can have that $\hat{\eta}_{k,t}=\hat{\eta}_{k,ite(k,t)}$, where $k\in[K],t\in[T]$.
@@
It’s easy to verify that $\hat{\eta}_{k_{t},t}=\eta_{next(k_{t},t)}=\eta_{t}$, where $t\in[T]$.
@@ C.3.3 Convergence Analysis for OrMo-DA
The second equation above holds because $\lceil\frac{t^{\prime}}{K}\rceil>\lceil\frac{t^{\prime}-1}{K}\rceil=I_{t^{\prime}}$ when $K\mid(t^{\prime}-1)$, $I_{t^{\prime}+1}=\lceil\frac{t^{\prime}}{K}\rceil$ and $\hat{\eta}_{k_{t^{\prime}},ite(k_{t^{\prime}},t^{\prime})}=\eta_{next(k_{t^{\prime}},ite(k_{t^{\prime}},t^{\prime}))}=\eta_{t^{\prime}}$. The last equation above holds because $\lceil\frac{t^{\prime}}{K}\rceil=\lceil\frac{t^{\prime}-1}{K}\rceil+1$.
@@
$~{}~{}~{}~{}\left.+\sum_{s=1}^{\lfloor\frac{t+K-2}{K}\rfloor}\sum_{j=(s-1)K+1}^{\min\{sK,t-1\}}\beta^{\lfloor\frac{t+K-2}{K}\rfloor-s}{\left(\hat{\eta}_{k_{j-1},j}\right)}^{2}\mathbb{E}\left\|\nabla F({\bf w}_{j})\right\|^{2}\right].$
@@
$\mathbb{E}\left\|\hat{{\bf u}}_{t}\right\|^{2}$
$\leq 2\mathbb{E}\left\|\beta^{\lfloor\frac{t+K-2}{K}\rfloor}\left(\sum_{k\in[K]}\hat{\eta}_{k,0}\nabla F({\bf w}_{0})\right)+\sum_{s=1}^{\lfloor\frac{t+K-2}{K}\rfloor}\beta^{\lfloor\frac{t+K-2}{K}\rfloor-s}\left(\sum_{j=(s-1)K+1}^{\min\{sK,t-1\}}\hat{\eta}_{k_{j-1},j}\nabla F({\bf w}_{j})\right)\right\|^{2}$
@@
$\leq\frac{2K}{1-\beta}\left[\sum_{k\in[K]}\beta^{\lfloor\frac{t+K-2}{K}\rfloor}{\left(\hat{\eta}_{k,0}\right)}^{2}\left\|\nabla F({\bf w}_{0})\right\|^{2}+\sum_{s=1}^{\lfloor\frac{t+K-2}{K}\rfloor}\sum_{j=(s-1)K+1}^{\min\{sK,t-1\}}\beta^{\lfloor\frac{t+K-2}{K}\rfloor-s}{\left(\hat{\eta}_{k_{j-1},j}\right)}^{2}\mathbb{E}\left\|\nabla F({\bf w}_{j})\right\|^{2}\right]$
$~{}~{}~{}~{}+2\sum_{k\in[K]}\beta^{2\lfloor\frac{t+K-2}{K}\rfloor}{\left(\hat{\eta}_{k,0}\right)}^{2}\sigma^{2}+2\sum_{s=1}^{\lfloor\frac{t+K-2}{K}\rfloor}\sum_{j=(s-1)K+1}^{\min\{sK,t-1\}}\beta^{2\lfloor\frac{t+K-2}{K}\rfloor-2s}{\left(\hat{\eta}_{k_{j-1},j}\right)}^{2}\sigma^{2}.$
@@
$\leq\frac{2\beta^{2}K\eta}{{(1-\beta)}^{3}}\sum_{t=1}^{T-1}\sum_{k\in[K]}\beta^{\lfloor\frac{t+K-2}{K}\rfloor}{\left(\hat{\eta}_{k,0}\right)}^{2}\left\|\nabla F({\bf w}_{0})\right\|^{2}$
$~{}~{}~{}~{}+\frac{2\beta^{2}K\eta}{{(1-\beta)}^{3}}\sum_{t=1}^{T-1}\sum_{s=1}^{\lfloor\frac{t+K-2}{K}\rfloor}\sum_{j=(s-1)K+1}^{\min\{sK,t-1\}}\beta^{\lfloor\frac{t+K-2}{K}\rfloor-s}{\left(\hat{\eta}_{k_{j-1},j}\right)}^{2}\mathbb{E}\left\|\nabla F({\bf w}_{j})\right\|^{2}$
$~{}~{}~{}~{}+\frac{2\beta^{2}\eta\sigma^{2}}{{(1-\beta)}^{2}}\sum_{t=1}^{T-1}\left[\sum_{k\in[K]}\beta^{2\lfloor\frac{t+K-2}{K}\rfloor}{\left(\hat{\eta}_{k,0}\right)}^{2}+\sum_{s=1}^{\lfloor\frac{t+K-2}{K}\rfloor}\sum_{j=(s-1)K+1}^{\min\{sK,t-1\}}\beta^{2\lfloor\frac{t+K-2}{K}\rfloor-2s}{\left(\hat{\eta}_{k_{j-1},j}\right)}^{2}\right]$
@@
$\sum_{t=1}^{T-1}\mathbb{E}\left(\hat{\eta}_{k_{t-1},t}\left\|\hat{{\bf w}}_{t}-{\bf w}_{t}\right\|^{2}\right)$ $\leq\frac{\eta K}{2L{(1-\beta)}^{2}}\left[\sum_{t=1}^{T-1}\hat{\eta}_{k_{t-1},t}\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}+\left(\sum_{k\in[K]}\hat{\eta}_{k,0}\right)\left\|\nabla F({\bf w}_{0})\right\|^{2}\right]$
@@
$\leq\frac{2\eta}{{(1-\beta)}^{2}}\sum_{j=1}^{T-2}{\left(\hat{\eta}_{k_{j-1},j}\right)}^{2}\left(\sigma^{2}+K\mathbb{E}\left\|\nabla F({\bf w}_{j})\right\|^{2}\right)\left(next(k_{j-1},j)-j\right)$
@@
$\mathbb{E}\left\|\nabla F(\bar{{\bf w}}_{T})\right\|^{2}\leq\mathcal{O}(\sqrt{\frac{L\sigma^{2}}{T}}+\frac{KL}{T}),$
where $\Delta=F({\bf w}_{0})-F^{*}$ and $\bar{{\bf w}}_{T}$ is randomly chosen from $\{{\bf w}_{0},{\bf w}_{1},\cdots,{\bf w}_{T-1}\}$ according to a probability distribution which is related to the delay-adaptive learning rates as shown in (17).
@@
According to Lemma 9, we can get that
@@
$\leq\sum_{t=1}^{T-1}\left[\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}\right]+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}$
@@
$\leq\sum_{t=1}^{T-1}\left[\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}\right]+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}$
@@
$\leq\sum_{t=1}^{T-1}\left[\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}\right]+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}$
@@
$\leq\sum_{t=1}^{T-1}\left[\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}\right]+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}$
@@
$\leq\sum_{t=1}^{T-1}\left[\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}\right]+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}$
@@
$\leq F({\bf w}_{0})+\left[\frac{L{\left(\sum_{k\in[K]}\hat{\eta}_{k,0}\right)}^{2}}{2{(1-\beta)}^{2}}-\frac{\sum_{k\in[K]}\hat{\eta}_{k,0}}{1-\beta}\right]\left\|\nabla F({\bf w}_{0})\right\|^{2}+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{k\in[K]}{\left(\hat{\eta}_{k,0}\right)}^{2}$
@@
$\leq\sum_{t=1}^{T-1}\left[\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{2(1-\beta)}\right)\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}\right]+\frac{L\sigma^{2}}{2{(1-\beta)}^{2}}\sum_{t=1}^{T-1}{(\hat{\eta}_{k_{t-1},t})}^{2}$
@@
$\leq\sum_{t=1}^{T-1}\left[\left(\frac{L{(\hat{\eta}_{k_{t-1},t})}^{2}}{2{(1-\beta)}^{2}}-\frac{\hat{\eta}_{k_{t-1},t}}{4(1-\beta)}\right)\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}\right]+\frac{2\eta L\sigma^{2}}{{(1-\beta)}^{3}}\left(\sum_{t=1}^{T-1}\hat{\eta}_{k_{t-1},t}+\sum_{k\in[K]}\hat{\eta}_{k,0}\right)$
@@
$\mathbb{E}F(\hat{{\bf y}}_{T})-F({\bf w}_{0})$ $\leq-\frac{1}{8\left(1-\beta\right)}\left[\sum_{t=1}^{T-1}\hat{\eta}_{k_{t-1},t}\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}+\sum_{k\in[K]}\hat{\eta}_{k,0}\left\|\nabla F({\bf w}_{0})\right\|^{2}\right]$
@@
$\frac{1}{\sum_{t=0}^{T-1}\bar{\eta}_{t}}\mathbb{E}\left(\sum_{t=0}^{T-1}\bar{\eta}_{t}\left\|\nabla F({\bf w}_{t})\right\|^{2}\right)\leq\frac{8\left(1-\beta\right)\left(F({\bf w}_{0})-F^{*}\right)}{\sum_{t=0}^{T-1}\bar{\eta}_{t}}+\frac{16\eta L\sigma^{2}}{{(1-\beta)}^{2}},$
where $\bar{\eta}_{0}=\sum_{k\in[K]}\hat{\eta}_{k,0}$ and $\bar{\eta}_{t}=\hat{\eta}_{k_{t-1},t}(t\geq 1)$.
Then, we analyse the lower bound of $\sum_{t=0}^{T-1}\bar{\eta}_{t}$.
@@
$\hat{\tau}_{k,t+1}=t+1-ite(k,t+1)=\left\{\begin{aligned} &t-ite(k,t)+1=\hat{\tau}_{k,t}+1&k\neq k_{t},\\ &0=\hat{\tau}_{k,t}-\tau_{t}&k=k_{t}.\end{aligned}\right.$
Since $\sum_{k\in[K]}\hat{\tau}_{k,0}=0$, we have $\sum_{k\in[K]}\hat{\tau}_{k,T}+\sum_{t=0}^{T-1}\tau_{t}=(K-1)T$. Moreover, $\hat{\tau}_{k_{T-1},T}=T-ite(k_{T-1},T)=0$. We get that $\sum_{k\in[K],k\neq k_{T-1}}\hat{\tau}_{k,T}+\sum_{t=0}^{T-1}\tau_{t}=(K-1)T$. Thus, at least $\frac{T}{2}$ delays are smaller than $2K$. $\sum_{t=0}^{T}\bar{\eta}_{t}=\sum_{t=1}^{T-1}\hat{\eta}_{k_{t-1},t}+\sum_{k\in[K]}\hat{\eta}_{k,0}\geq\frac{T\eta}{2}$.
$\frac{1}{\sum_{t=0}^{T-1}\bar{\eta}_{t}}\sum_{t=0}^{T-1}\bar{\eta}_{t}\mathbb{E}\left\|\nabla F({\bf w}_{t})\right\|^{2}\leq\frac{16(1-\beta)}{T\eta}\left(F({\bf w}_{0})-F^{*}\right)+\frac{16\eta L\sigma^{2}}{{(1-\beta)}^{2}}.$
If we choose an output $\bar{{\bf w}}_{T}$ from $\{{\bf w}_{0},{\bf w}_{1},\cdots,{\bf w}_{T-1}\}$ according to
${\mathbb{P}}(\bar{{\bf w}}_{T}={\bf w}_{t})\propto\bar{\eta}_{t},$ (17)
where $t\in[T]$ and let $\eta=\min\{\frac{(1-\beta)^{2}}{8KL},\sqrt{\frac{(1-\beta)^{3}\left(F({\bf w}_{0})-F^{*}\right)}{TL\sigma^{2}}}\}$, we have that
$\mathbb{E}\left\|\nabla F(\bar{{\bf w}}_{T})\right\|^{2}\leq\mathcal{O}\left(\sqrt{\frac{L\sigma^{2}}{T}}+\frac{KL}{T}\right).$